OpenStudy (anonymous):

Find the max/min values for P(x) = -2x^2 +8x-1

OpenStudy (anonymous):

on a graphing calculator it reads that 7 is the local maximum (y=7)

OpenStudy (raden):

*The first way, find the x-vertec by use formula x=-b/2a = -8/2(-2) = 2 to find the max value, just plug x=2 to P(x) so, P(2) = -2(2)^2+8(2)-1=7 *The 2nd way, use the formula y=-D/4a (with D=b^2-4ac) y_max = -(8^2-4(-2)(-1))/4(-2) = -(64-8)/-8 = -56/-8 = 7

OpenStudy (ash2326):

@oheneba do you know calculus?

OpenStudy (anonymous):

To an extent

OpenStudy (zzr0ck3r):

you dont even need calculus, just use the vertex formula you learn in algebra. you know it is a parabola that opens down, so you will have a max value. it will be the y component of the vertex

OpenStudy (anonymous):

Thanks